# ChatReID: Open-ended Interactive Person Retrieval via Hierarchical Progressive Tuning for Vision Language Models

### v1

@@ Abstract
Person re-identification (Re-ID) is a critical task in human-centric intelligent systems, enabling consistent identification of individuals across different camera views using multi-modal query information. Recent studies have successfully integrated LVLMs with person Re-ID, yielding promising results. However, existing LVLM-based methods face several limitations. They rely on extracting textual embeddings from fixed templates, which are used either as intermediate features for image representation or for prompt tuning in domain-specific tasks. Furthermore, they are unable to adopt the VQA inference format, significantly restricting their broader applicability. In this paper, we propose a novel, versatile, one-for-all person Re-ID framework, ChatReID. Our approach introduces a Hierarchical Progressive Tuning (HPT) strategy, which ensures fine-grained identity-level retrieval by progressively refining the model’s ability to distinguish pedestrian identities. Extensive experiments demonstrate that our approach outperforms SOTA methods across ten benchmarks in four different Re-ID settings, offering enhanced flexibility and user-friendliness. ChatReID provides a scalable, practical solution for real-world person Re-ID applications, enabling effective multi-modal interaction and fine-grained identity discrimination.
@@ 1 Introduction
Figure 1: Illustration of ChatReID’s capabilities across different person re-identification tasks. (a) shows the Standard Person Re-ID task, (b) shows Visible-infrared Person Re-ID task, (c) shows Text-to-image Re-ID task, and (d) shows Adaptive Multi-Modal Person Retrieval task.
Person re-identification (Re-ID) is a crucial task aimed at identifying individuals across different camera views using multi-modal query information. It plays a vital role in human-centric intelligent systems with broad practical applications in crowd management, smart city infrastructure, and security analytics. In recent years, large vision-language models (LVLMs) have made substantial advancements, showcasing remarkable performance by leveraging extensive world knowledge and an enhanced ability to interpret human instructions. For example, Qwen2-VL [37] introduces Naive Dynamic Resolution and Multimodal Rotary Position Embedding (M-RoPE) to improve cross-modal fusion, significantly enhancing capabilities in image and video understanding. Although some of them also have the ability to distinguish different images, they can hardly perceive fine-grained information to solve person Re-ID.
Currently, many LVLM-based approaches for person Re-ID can be broadly categorized into three main types. 1) Cross-Modal Pedestrian Features Learning: These approaches leverage the alignment capabilities of LVLMs between text and image embeddings by incorporating template language descriptions, such as “A photo of []” or “A person with [].” 2) Text-to-Image Person Re-ID: These approaches focus on retrieving target identity images using textual descriptions as queries, leveraging the alignment capabilities of LVLMs between text and image features. 3) Unified Multi-Task Framework: These methods exploit the multimodal reasoning capabilities of LVLMs to integrate different Re-ID tasks within a unified framework, facilitating a cohesive Re-ID system. However, all of them preclude the adoption of the widely recognized LVLM-based VQA inference format. It is difficult to extend these models to common application scenarios.
In this paper, we concentrate on exploring complex-content reasoning and multi-modal retrieval capabilities of LVLMs to tackle the aforementioned limitations. We develop a versatile ‘one-for-all’ person Re-ID framework, ChatReID, to interactively ask the machine to help re-identify a person given arbitrary, freeform, and necessary descriptions or clues. An illustration of some functionalities is shown in Figure 1, which demonstrates the capability to perform VQA inference format person Re-ID as well as adaptive multi-modal person retrieval. Nevertheless, achieving this goal presents significant challenges without large-scale pre-training. This complexity stems primarily from two factors: 1) Fine-Grained Identity-Level Retrieval: Unlike general image retrieval tasks, person Re-ID demands fine-grained, identity-level differentiation. While existing LVLMs can easily distinguish between broad categories (e.g., cats and dogs), they lack the specialized pre-training required to recognize individual identities based on nuanced visual or textual cues. 2) Diverse Re-ID Task Settings: Different Re-ID tasks focus on unique sets of visual features, which complicates integrating them effectively within a single model and evaluating performance using a unified VQA inference format.
To achieve this goal, we design a Hierarchical Progressive Tuning (HPT) strategy, which can be divided into three training stages: incremental pre-training for the ReID-specific vision encoder, scaling up pedestrian discrimination capability via multi-task joint training, and enhancing the pedestrian retrieval capability via post-training the decoder. Through HPT, the model’s inherent matching and reasoning capabilities are gradually enhanced to achieve fine-grained identity discrimination. In Stage 1, we perform incremental pre-training of the vision encoder to help the vision encoder understand identity-specific features of pedestrians. In Stage 2, we conduct joint training with seven tasks, enabling LVLMs to acquire an initial capability for person identity matching based on images, textual descriptions, and pedestrian attributes. In Stage 3, we design instruction fine-tuning for three specific application scenarios to enhance the practicality of ChatReID and deepen the model’s understanding of the person Re-ID task objective. As shown in Figure 2, extensive experiments demonstrate that our method achieves SOTA results across ten benchmarks in four different person Re-ID settings.
Figure 2: Performance across ten benchmarks in four different person Re-ID settings. This demonstrates ChatReID’s robustness and effectiveness across diverse Re-ID challenges.
Our paradigm has several notable characteristics:
$\bullet$ Leveraging LVLM Retrieval Capabilities. Through a Hierarchical Progressive Tuning strategy, our model fully develops the fine-grained identity discrimination capabilities of LVLMs, closely aligned with real-world application requirements.
$\bullet$ User-Friendly Design. By utilizing a VQA inference format, ChatReID makes complex person re-identification tasks accessible, enabling users to handle these tasks effectively without extensive technical expertise.
$\bullet$ Enhanced Flexibility. ChatReID can not only support traditional Re-ID tasks but also allow users to freely combine input modalities or adjust the sequence of inputs, resulting in a more flexible and accurate person Re-ID system suited to practical applications.
@@ 2.1 Traditional Person Re-ID
Person re-identification (Re-ID) is a fundamental task in computer vision, which aims to match the same individual across different camera views based on visual features. Recent studies in person Re-ID carefully designed settings and developed models to tackle every specific scenario. Standard person Re-ID [59, 60, 26, 34, 14, 52], which aims to match individuals across cameras based on visual features. These methods distinguish pedestrian identities based on body posture and appearance. Cloth-changing Re-ID (CC Re-ID) [27, 4, 20, 15, 10] is a more challenging variant where individuals change their clothing between camera views. It assists the model in extracting non-clothing information for identity determination. CSSC [38] introduces a framework that effectively leverages abundant semantics within pedestrian images to extract identity features. Visible-infrared person ReID (VI-ReID) methods [7, 39, 17] extract pedestrian features under low-light environments. DDAG [49] improves performance by leveraging intra-modality and cross-modality contextual cues to enhance feature discriminability and robustness to noise. Text-to-image Re-ID [29, 11] aims to identify pedestrians based on textual descriptions. It requires the model to understand and align linguistic descriptions with visual attributes. Zhao et al. [57] proposes a novel method to model multi-modal uncertainty and semantic alignment using Gaussian distributions and a cross-modal circle loss. However, different settings within person Re-ID focus on distinct visual features, making it difficult to effectively integrate these settings into a single model. Consequently, we intend to develop a versatile ‘one-for-all’ framework to interactively ask the machine to help with the person retrieval task.
2.2 LVLM-driven Person Re-ID
@@
Large Vision-Language Models (LVLMs) [2, 37, 48, 24] have garnered significant attention in the AI community due to their impressive generalization capabilities. Recent studies have started investigating the incorporation of LVLMs into the person Re-ID paradigm. Tan et al.[35] and Yang et al.[46] primarily focuses on the text-to-image person Re-ID task. The former uses multi-modal large language models (MLLMs) to caption images according to various templates, thereby addressing issues related to the quantity and quality of textual descriptions. The latter propose a common instruction template and use features computed by MLLMs to train person Re-ID models. Instruct-ReID [13] is the first work that unifies multiple person Re-ID settings within a single model, generating task-specific instructions and combining instruction encodings with visual encodings for re-identification training. Despite significant progress in integrating LVLMs into person Re-ID, existing methods face key limitations. Firstly, they fail to fully utilize LVLMs’ perception and instruction-following abilities. Secondly, many approaches rely on rigid, template-based textual descriptions, limiting adaptability and scalability. Lastly, while some methods unify different Re-ID settings, their flexibility remains constrained, making it difficult to apply them to common scenarios. In this paper, we present a versatile ‘one-for-all’ Re-ID framework that leverages LVLMs for interactive, freeform person Re-ID.
@@ 3 Methodology
In this paper, we propose an innovative person re-identification paradigm called ChatReID, which achieves an open-ended, interactive Re-ID application system. The subsequent sections present a comprehensive overview of the proposed methodological framework, the hierarchical progressive tuning strategy, and the data engine.
Table 1: Statistics of Datasets Used in Each Training and Inference Stage. For image-based datasets, we present the number of images and IDs. For text-to-image datasets, we show the number of images and captions.
Phase Datasets #Images #IDs/Captions
Stage 1 LUPerson-NL [8] 10,683,716 433,997
OmniReID [13] 4,973,044 328,604
Stage 2 DeepChange [43] 187,000 1,121
COCAS-Real2 [51] 21,319 101
LLCM [54] 46,767 1,064
LaST [30] 228,000 10,862
Flickr 30k [50] 31,783 158,915
UFineBench [64] 33,652 90,351
SYNTH-PEDES[63] 4,791,771 312,321
Stage 3 / Infer Market1501 [58] 32,668 $1,501$
MSMT17 [40] 126,441 4,101
CUHK03 [22] 14,096 1,467
PRCC [45] 33698 221
LTCC [27] 17,138 152
SYSU-MM01 [41] 45,863 491
RegDB [25] 8,240 412
CUHK-PEDES [21] 43,264 86,528
ICFG-PEDES [5] 54,522 54,522
RSTPReid [61] 20,505 41,010
Framework. In this section, we detail the technical architecture of ChatReID. The ChatReID framework, illustrated in Figure 3, integrates the structure of Qwen2-VL [37] due to its Naive Dynamic Resolution architecture, which effectively addresses the challenges posed by variable image resolutions in person Re-ID tasks. The hierarchical tuning process consists of three stages: Incremental Pre-training aims to help the vision encoder capture identity-specific features of pedestrians, enabling it to focus on fine-grained details relevant to person Re-ID. Both the encoder and decoder are initialized with a pre-trained LVLM to leverage existing language-vision knowledge. Multi-task Joint Training develops foundational skills in person identity matching based on images, textual descriptions, and pedestrian attributes. Only the encoder developed in Stage 1 is retained. Decoder Post-training enhances ChatReID’s understanding of Re-ID objectives and improves its instruction-following capabilities for practical applications. Here, we build upon the model from Stage 2 and freeze the encoder.
Figure 3: Overview of ChatReID’s three stage hierarchical training framework.
Training Datasets. For each training stage, we collect and organize a substantial number of person Re-ID datasets to construct the instruction tuning datasets used for training. The statistics of the datasets used in each stage are shown in Table 1. In Stage 1, we collect and organize two large-scale datasets, LUPerson-NL [8] and OmniReID [13], to construct a person image caption instruction dataset with over 5 million samples. In Stage 2, we expand our dataset collection to include a range of public datasets from various Re-ID settings. These include DeepChange [43], COCAS-Real2 [51], and LaST [30] for standard person Re-ID and CC Re-ID; LLCM [54] for VI-ReID; and UFineBench [64] and SYNTH-PEDES [63] for text-to-image Re-ID, totaling 5,340,292 images. Using these datasets, we constructed a joint training instruction dataset comprising 3 million samples across seven training tasks. In Stage 3, we gather ten widely used benchmarks across four Re-ID settings to create a comprehensive instruction dataset with 500,000 samples and three specific training objectives. These benchmarks include Market1501 [58], MSMT17 [40], and CUHK03 [22] for standard person Re-ID; PRCC [45] and LTCC [27] for CC Re-ID; SYSU-MM01 [41] and RegDB [25] for VI-ReID; and CUHK-PEDES [21], ICFG-PEDES [5], and RSTPReid [61] for text-to-image Re-ID. During the training phase, we utilized only the training subsets of these benchmarks, while the inference phase involved evaluation using their respective test sets.
3.1 Incremental Pre-training
3.1.1 Data Engine
In this stage, we utilize over 5M image-text instruction pairs. We begin by collecting and processing two large-scale person Re-ID datasets, where many images initially exhibited issues like blocking, blurring, and the presence of multiple individuals, rendering them unsuitable for pre-training. To address this, we performed post-processing to reduce noise and enhance data quality. The data construction process involved several key steps: Image Filtering by Quality. We filter images based on size and resolution to address blurriness and improve overall quality. Image Filtering by Attributes. We employ GPT-4o [1] to predict pedestrian attributes. If the model failed to recognize the pedestrian’s clothing. These are marked as occluded and removed from pre-training. Image Captioning. Using GPT-4o, we generated high-quality, diverse textual descriptions for each image. Multiple descriptions are created for each image with varying syntactic structures to enhance variability and richness.
3.1.2 The Vision Encoder pre-training
We conduct full-parameter training on both the vision encoder and language decoder of Qwen2-VL 2b. In Stage 1, full-parameter training enhances the encoder’s focus on pedestrian features, while training the decoder ensures that captions accurately reflect pedestrian attributes. The patch merger further guides the encoder’s training effectively. We standardize images with resolutions below 256x128, resizing them to 256x128 while leaving higher-resolution images unchanged. This adjustment leverages our experience with pedestrian tasks, and moderate image compression enhances training efficiency.
3.2 Multi-task Joint Training
After completing Stage 1, we obtain an encoder capable of focusing on the visual features specific to pedestrians. However, at this stage, the encoder-decoder still lacks the ability to distinguish individual identities. To address this, we implement a multi-task joint training strategy and an enhanced data engine to incorporate comprehensive pedestrian identification knowledge into ChatReID.
@@ 3.2 Multi-task Joint Training
Figure 4: Multi-Task Design Diagram for joint training in Stage2. Concretely, we conduct seven distinct matching and retrieval tasks between text and image modalities, encouraging LVLMS to acquire an initial capability for person identity matching based on images, textual descriptions, and pedestrian attributes.
3.2.1 Multi-task Design
Multi-Task Design Diagram as shown in Figure 4. We structure the pedestrian discrimination task into three objectives: image-to-image retrieval, text-to-image retrieval, and attribute-based retrieval. Image-to-image retrieval includes standard person Re-ID, CC Re-ID, and VI-ReID. Text-to-image retrieval is used to match individuals based on textual descriptions. Attribute-Based Retrieval focuses on distinguishing individuals based on specific pedestrian attributes, which play a crucial role in person retrieval. Additionally, our multi-task design incorporates varying levels of input complexity: one-to-one pairing, one-to-many retrieval, and many-to-many retrieval. This diverse range of task complexities is essential for balancing training difficulty and achieving robust results.
Image-to-image One-to-One Matching. We sample two images from the dataset and query whether the pedestrians shown in both images belong to the same identity. To ensure an adequate number of positive examples for effective learning, we set a probability threshold for the sampled images to belong to the same identity.
Image-to-image Many-to-Many Retrieval. We sample an image as the query image and a set of $N$ images as the gallery, where $N$ includes $n$ images that share the same identity as the query image. The task is to assess whether the model can accurately locate all images corresponding to the query identity. Both $N$ and $n$ vary dynamically.
Text-to-image One-to-One Matching. We sample an image and a text description from the dataset and query if they are a matching pair. To ensure an adequate number of positive examples for effective learning, we set a probability threshold for the samples to belong to the same identity.
Text-to-image Many-to-One Retrieval. We sample an image and multiple text descriptions from the dataset, then query which caption accurately describes the query image.
Text-to-image One-to-Many Retrieval. We sample multiple images and a single text description from the dataset, then query which image accurately matches the description.
Attribute-level Commonality and Uniqueness Comparison. We sample two images and query the model to explain the similarities and distinctions in the identifiable attributes of the pedestrians shown. This approach enhances the model’s comprehension of pedestrian attributes. In this case, ground truth data is derived from a text-to-image dataset with attribute annotations. We manually extract shared and unique attributes from the two images to serve as supervisory information during training.
Attribute Annotations Prediction. We sample an image and query the model to predict its attribute annotations (e.g., gender, clothing color, etc.).
Instead of manual text query inputs, we create a series of multi-turn dialogue data$(X_{1}^{q},X_{1}^{a},\cdots,X_{T}^{q},X_{T}^{a})$, with T representing the complete count of dialogue exchange, addressing the divergence in multi-task interpretation between humans and large language models. For each individual task, the query format is altered, and only one instance is provided in the text.
3.2.2 Data Engine

### v2

@@ Abstract
Person re-identification (Re-ID) is a crucial task in computer vision, aiming to recognize individuals across non-overlapping camera views. While recent advanced vision-language models (VLMs) excel in logical reasoning and multi-task generalization, their applications in Re-ID tasks remain limited. They either struggle to perform accurate matching based on identity-relevant features or assist image-dominated branches as auxiliary semantics. In this paper, we propose a novel framework ChatReID, that shifts the focus towards a text-side-dominated retrieval paradigm, enabling flexible and interactive re-identification. To integrate the reasoning abilities of language models into Re-ID pipelines, We first present a large-scale instruction dataset, which contains more than 8 million prompts to promote the model fine-tuning. Next. we introduce a hierarchical progressive tuning strategy, which endows Re-ID ability through three stages of tuning, i.e., from person attribute understanding to fine-grained image retrieval and to multi-modal task reasoning. Extensive experiments across ten popular benchmarks demonstrate that ChatReID outperforms existing methods, achieving state-of-the-art performance in all Re-ID tasks. More experiments demonstrate that ChatReID not only has the ability to recognize fine-grained details but also to integrate them into a coherent reasoning process.
@@ 1 Introduction
Figure 1: Performance comparisons across ten benchmarks on four person Re-ID tasks show that our ChatReID significantly outperforms previous state-of-the-art methods, demonstrating its superior robustness and effectiveness.
Figure 2: (a) and (b) show the results of standard and cloth-changing(CC) Re-ID using GPT-4o and Qwen2-VL without tuning. (c)–(g) illustrate ChatReID’s capabilities across various person Re-ID tasks. Red dialogue boxes indicate incorrect responses, while green dialogue boxes indicate correct responses. Best viewed in color and zoom in.
Person re-identification (Re-ID) is a fundamental task in computer vision that aims to recognize individuals across non-overlapping camera views [55, 49, 14]. It plays a crucial role in intelligent surveillance systems, facilitating applications in suspect tracking, crowd management, and access control. In the past decade, person Re-ID has achieved remarkable progress with the development of deep learning techniques [7, 50, 4], significantly relieving challenges of pose variants, illumination and occlusions.
Recently, the emergence of vision-language models (VLMs) [1, 39, 46, 26] present new Re-ID paradigms by integrating textual descriptions with visual representations. They leverage the ability of image-text alignment to enrich feature representations, further promoting the performance and application range of person Re-ID. Depending on the usage of text information, VLM-based methods can be broadly grouped into three categories. (1) Auxiliary supervision using text [23, 47]: they employ text embedding to additionally supervise the feature learning or to integrate with image features. (2) Text-based person retrieval [37, 45]: they adopt textual descriptions as queries to retrieve images with the matched identity. (3) Text prompts for unified Re-ID [12, 13]: they use text prompts or language instructions to manipulate a unified model handling different Re-ID tasks. Despite the impressive effect, these methods still follow the traditional image-side dominated Re-ID paradigm, i.e., extracting features with feature extractors (or encoders) and then sorting them by calculating feature similarity.
Essentially, person Re-ID is a complex reasoning process that requires analyzing a person’s appearance and biological information from an image, then repeatedly comparing different and similar points in a pair of images to determine whether their identities match. However, existing VLMs, such as Qwen2-VL [39] and GPT-4o [18], posses strong vision-language reasoning abilities in programming and communication, yet their potential in person Re-ID tasks remains largely unstudied. This raises a natural question: can we leverage the advanced reasoning capabilities of VLMs to perform text-side dominated person Re-ID? Intuitively, a straightforward practice is to retain the text decoder of VLMs in the pipeline and ask it to provide the index of the matched images or the similarity score between the given query and gallery images. However, as depicted in Fig. 2(a) and (b), GPT-4o focuses only on appearance features, leading to incorrect matches, while Qwen2-VL produces almost illogical results.
We argue that successful person Re-ID requires not only recognizing fine-grained details but also integrating them into a coherent reasoning process. Concretely, (1) person identity is inherently an intra-class semantics, meaning that person Re-ID relies on fine-grained biological or identity-related features. While existing VLMs can easily distinguish between broad categories (i.e., cats v.s. dogs) or major attributes (i.e., clothing colors), they lack the specialized tuning needed to identify individuals based on subtle visual or textual clues. (2) In addition, person Re-ID is an open-set task, where the correct matches can vary depending on the given query or task context. Unlike approaches that use text prompts as conditions, the text-side dominated Re-ID framework requires inferring the desirable similarities through a deep and comprehensive understanding of both textual descriptions and visual content.
In this paper, we explore complex-content reasoning and multi-modal retrieval capabilities of VLMs and introduce ChatReID, a new perspective of using the text decoder for person Re-ID. Our ChatReID is a versatile ‘one-for-all’ framework to interactively ask the model to assist in re-identifying a person given arbitrary, freeform, and necessary descriptions or clues, as shown in Fig. 2(c)-(g). To overcome the aforementioned challenges, we first propose a hierarchical progressive tuning (HPT) strategy, which consists of three stages. The first stage helps the model understand the attributes or semantics of a person. In the second stage, the model is guided to learn the ability of image-to-image, image-to-attribute, and text-to-image fine-grained retrieval. The third stage deepens the logical reasoning ability of the model between application scenario descriptions and inputs. Secondly, we create a large-scale instruction dataset with more than 8 million prompts, to promote the tuning of three progressive stages.
Contributions. We summarize contributions as follows:
  • We propose ChatReID, a novel framework for person Re-ID that introduces a text-side-dominated retrieval paradigm. ChatReID enables a flexible and interactive retrieval process, enhancing stronger generalization ability.
  • We present a large-scale instruction dataset and a hierarchical progressive tuning strategy, which endows Re-ID ability through three stages of tuning, i.e., from person attribute understanding to fine-grained image retrieval and to multi-modal task reasoning.
  • Extensive experiments on ten widely used benchmarks across four different person Re-ID tasks to evaluate the effectiveness of our model. ChatReID achieves state-of-the-art performance in all experiments, outperforming existing methods by a significant margin.
@@ 2.1 Traditional Person Re-ID
Person re-identification (Re-ID) is a fundamental task in computer vision, which aims to match the same individual across different camera views based on visual features. Recent studies in person Re-ID carefully designed settings and developed models to tackle every specific scenario. Standard person Re-ID [55, 56, 28, 36, 14, 49], which aims to match individuals across cameras based on visual features. These methods distinguish pedestrian identities based on body posture and appearance. Cloth-changing Re-ID (CC Re-ID) [29, 3, 21, 15, 9] is a more challenging variant where individuals change their clothing between camera views. It assists the model in extracting non-clothing information for identity determination. CSSC [40] introduces a framework that leverages abundant semantics within pedestrian images to extract identity features. Visible-infrared person ReID (VI-ReID) methods [6, 41, 17] extract pedestrian features under low-light environments. DDAG [48] improves performance by leveraging intra-modality and cross-modality contextual cues to enhance feature discriminability and robustness to noise. Text-to-image Re-ID [32, 10] aims to identify pedestrians based on textual descriptions. It requires the model to understand and align linguistic descriptions with visual attributes. Zhao et al. [54] proposes a novel method to model multi-modal uncertainty and semantic alignment using Gaussian distributions and a cross-modal circle loss. However, different settings within person Re-ID focus on distinct visual features, making it difficult to effectively integrate these settings into a single model. Consequently, we intend to develop a versatile ‘one-for-all’ framework to interactively ask the machine to help with the person retrieval task.
2.2 VLM-driven Person Re-ID
@@
Vision-language models (VLMs) [1, 39, 46, 26] have garnered significant attention in the AI community due to their impressive generalization capabilities. Recent studies have started investigating the incorporation of VLMs into the person Re-ID paradigm. Tan et al.[37] and Yang et al.[44] primarily focuses on the text-to-image person Re-ID task. The former uses multi-modal large language models (MLLMs) to caption images according to various templates, thereby addressing issues related to the quantity and quality of textual descriptions. The latter proposes a common instruction template and uses features computed by MLLMs to train person Re-ID models. Instruct-ReID [12] is the first work that unifies multiple person Re-ID settings within a single model, generating task-specific instructions and combining instruction encodings with visual encodings for Re-ID training. Despite significant progress in integrating VLMs into person Re-ID, existing methods face key limitations. Firstly, they fail to fully utilize VLMs’ perception and instruction-following abilities. Secondly, many approaches rely on rigid, template-based textual descriptions, limiting adaptability and scalability. Lastly, while some methods unify different Re-ID settings, their flexibility remains constrained, making it difficult to apply them to common scenarios. In this paper, we present a versatile ‘one-for-all’ Re-ID framework that leverages VLMs for interactive, freeform person Re-ID.
Figure 3: Overview of the ChatReID framework. (a) shows the schematic of ChatReID. (b) shows the three-stage HPT strategy.
@@ 3 Methodology
Our ChatReID is a text-side-dominated framework for person Re-ID. Different from traditional methods that calculates image feature distance as similarity, ChatReID interprets the task requirements from the input text description and performs similarity inference on the given person images accordingly. Finally, it outputs the matched person image in textual form.
Figure 3(a) illustrates the schematic of our ChatReID, which is primarily composed of a ViT encoder and a LLM decoder. In this section, we start by introducing how to fine-tune the encoder and decoder through our proposed hierarchical progressive tuning strategy (in 3.1). Next, we elaborate a large-scale customized instruction dataset built for fine-tuning (see 3.2). Lastly, we discuss the architecture details of our framework, along with the training and inference processes in Sec. 3.3.
3.1 Hierarchical Progressive Tuning
Person Re-ID is a fine-grained retrieval task where identity represents highly abstract semantic information, making it difficult to achieve accurate matches using existing VLMs directly, such as Qwen2-VL [39] and GPT-4o [18] shown in Fig. 2(a)-(b). However, VLMs excel in logical reasoning and multi-task generalization, offering great potential for Re-ID. To bridge this gap, as shown in Fig. 3(b), we introduce a hierarchical progressive fine-tuning strategy, which gradually enhances the model’s Re-ID capability through three stages: (1) person attribute understanding, (2) fine-grained image retrieval, and (3) multi-modal task reasoning.
3.1.1 Stage One: Person Attribtue Understanding
Considering the diversity of person images, the first stage focuses on improving the model’s ability to understand pedestrians by learning and extracting fine-grained attributes such as gender, clothing, and posture.
$\bullet$ Image captioning. Following previous studies [25, 22], we use image captioning as a foundational fine-tuning task. Specifically, the model is trained to generate detailed attribute descriptions based on a given person image and corresponding prompts. To further guide the model, we empirically incorporate several predefined key attributes into the prompt. This design not only encourages the model to capture more precise and informative attribute descriptions but also potentially strengthens its ability to differentiate identities. The prompt used is:
“In the {person image} provided, can you give a detailed description of the pedestrian, including their gender, age range, hair, type and color of clothing, footwear type and color, posture or gait, any patterns or accessories, and any distinguishing features?”
3.1.2 Stage Two: Fine-grained Image Retrieval
After completing the first stage of tuning, we obtain a vision encoder that can effectively capture fine-grained person attributes. To further promote its ability for person re-identification, we introduce several fine-tuning tasks in the second stage to guide the model in matching identity-related features with fine-grained information. Specifically, we structure tasks into three categories of retrieval tasks: image-to-image, text-to-text, and image-to-attribute.
Additionally, each type of retrieval task is designed with varying levels of complexity, including one-to-one, one-to-many, and many-to-many scenarios. This diverse range of task complexities is crucial for balancing training difficulty and improving the model’s robustness (see our discussions in Sec. 4.3). Figure 4 illustrates seven fine-tuning tasks, including examples of prompts and responses, involved in the second stage.
@@ 3.1.2 Stage Two: Fine-grained Image Retrieval
Figure 4: Multi-task Design Diagram for joint training in the stage 2. Concretely, we conduct seven distinct matching and retrieval tasks between text and image modalities, encouraging VLMS to acquire an initial capability for fine-grained image retrieval based on images, textual descriptions, and pedestrian attributes.
$\bullet$ Image-to-image Retrieval. This task has two levels of one-to-one and one-to-many. For the one-to-one level, we sample two images and ask the model to determine whether the people in these two images are the same person. To balance the training, we set a probability of 0.5 to sample images with positive answers. For the one-to-many level, we randomly select one image as the query and choose a set of $N$ images, where $N$ images include $n$ images that have the same identity. The model is required to identify all matched images. During training, values of both $N$ and $n$ vary to make the task more challenging and diverse.
$\bullet$ Text-to-image Retrieval. This task aims to identify the matched person image based on a text description, which has three levels of one-to-one, many-to-one, and one-to-many. Similar to the image-to-image retrieval task, the one-to-one level requires the model to determine if the image and the text description match or not (‘yes’ or ‘no’). And for the last two levels, both need to find the correctest answer from a set of options.
$\bullet$ Image-to-attribute Retrieval. This task is proposed to learn to distinguish individuals based on specific attributes, and we design two specific levels. For the first one, two images are sampled and the model is tasked with explaining the similarities and distinctions in attributes of two pedestrians. It is expected to further enhance the model’s understanding ability of person attributes. For the second, an image is randomly selected and fed into the model to generate its attribute descriptions or annotations (e.g., gender, clothing color, etc.). It is similar to the first stage, with the intention to deepen and consolidate this capability of the model in the current stage.
3.1.3 Stage Three: Multi-modal Task Reasoning
Stage Three enhances ChatReID’s rasoning capability of Re-ID objectives and improves its instruction-following capabilities for practical applications. After compressing the fine-grained pedestrian discrimination information via the above two stages, ChatReID has foundational skills in person identity matching. In this stage, we directly adopt the training objective of 4 person Re-ID tasks.
$\bullet$ Standard person Re-ID. Standard person Re-ID is a image-to-image retrieval task, where a query image and a set of gallery images are provided, and the goal is to identify the pedestrian in the gallery that matches the identity of the query, the prompt used is:
“Given the {person image} provided, identify which of the following images show the same pedestrian as in the first image.”
$\bullet$ CC person Re-ID. Based on the cloth-changing person Re-ID setting, the prompt used is:
“Given the {person image} provided, select which of the {gallery} shows the same person, who may be wearing different clothing.”
$\bullet$ VI person Re-ID. Based on the VI person Re-ID setting, the prompt used is:
“Using the given {person image}, choose the corresponding {gallery} that matches the same pedestrian, considering both infrared and visible light images for re-identification.”
$\bullet$ T2I person Re-ID. T2I person Re-ID is a text-to-image retrieval task, where a pedestrian image description is given, and the goal is to find the matching pedestrian image in the gallery based on the description, the prompt used is:
“Select the {person image} that best matches the following description of the pedestrian’s appearance: {text} ’̇’
To improve robustness and prevent bias from fixed-length galleries, we introduce variability in the gallery length, randomly adjusting it while ensuring at least one image matches the query identity.
3.2 Data Engine
For each training stage, we collect and organize a substantial number of person Re-ID datasets to construct the instruction tuning datasets used for training. We utilize a total of 19 open-source person Re-ID datasets. The complete list of datasets and corresponding statistical information can be found in the supplementary materials.
In Stage 1, we utilize over 5M image-text instruction pairs. To address the low-quality problem, we performed post-processing to reduce noise and enhance data quality. The data construction process involved several key steps: Image Filtering by Quality. We filter images based on size and resolution to address blurriness and improve overall quality. Image Filtering by Attributes. We employ GPT-4o [18] to predict pedestrian attributes. If the model failed to recognize the pedestrian’s clothing. These are marked as occluded and removed from pre-training. Image Captioning. Using GPT-4o [18], we generated high-quality, diverse textual descriptions for each image.